OpenStudy (anonymous):

Which of the following makes this a true equation? [] + 1 1/4 = 4 1/2 2 1/2 2 3/4 3 1/4 3 1/2

OpenStudy (aroub):

3 1/4

OpenStudy (anonymous):

Set the unknown to x giving you x + 1(1/4) = 4(1/2) Then simplify: x + 1x1/4 = 4x1/2 => x+1/4 = 4/2 Then subtract 1/4 from both sides => x + 1/4 - 1/4 = 4/2 - 1/4 Then find a common denominator => x = 8/4 - 1/4 Then simplify => x = 7/4 //That't gives you the answer. Proof: 7/4 + 1/4 = 4 1/2 => 8/4 = 4 1/2 => 8/4 = 4/2 => 2 = 2